# Joint Modelling of Astrophysical Systematics for Cosmology with LSST

### v2

@@ 4.4.3 Degeneracies: within the luminosity function parameter space
In analyzing the correlations and degeneracies between the Schechter luminosity function parameters themselves, particularly considering their tuning for the LSST Year 1 and Year 10 forecasts, we observe a nuanced interplay. For $\boldsymbol{\phi}_{0}^{*}$, the absence of correlation with $P$ in both Year 1 and Year 10 suggests that the normalization factor of the galaxy number density operates somewhat independently of its evolution parameter. In fact, $\phi_{0}^{*}$ has minimal degeneracy with any of the other luminosity function parameters, which is perhaps unsurprising since it serves only as an overall normalisation whereas the others all provide some redshift evolution effects. Parameter $\boldsymbol{P}$ exhibits a strong anti-correlation with $M_{0}^{*}$, $Q$, and $\alpha$ in both years, which can be interpreted as being due to the fact that this set of parameters all induce shifts in the means of the source redshift distributions. The strong positive correlation of $\boldsymbol{M_{0}^{*}}$ with both $Q$ and $\alpha$ in both years, tightening in Year 10, further highlights this effect. $\boldsymbol{Q}$ shows an evolving relationship with $\boldsymbol{\alpha}$, transitioning from a positive correlation in Year 1 to a slightly negative one in Year 10. We expect that this is due to the different fiducial values of the LF parameters, corresponding to a shift in how the rate of change in galaxy luminosity correlates with the abundance of fainter galaxies over time.
The complex interplay between luminosity function (LF) parameters, intrinsic alignment (IA) models, and cosmological parameters underscores the need for an even more extensive investigation. Notably, the parameter space of the LF can be reduced from five to three parameters, as defined in Eq. 4, streamlining our analysis without sacrificing depth or accuracy. Given the intricacies of how these elements influence each other, a comprehensive approach is necessary. A Markov Chain Monte Carlo method, with its capability to dispense with simplifying assumptions of Fisher forecasting (e.g., Gaussian likelihoods), presents a promising route. This approach allows for a thorough examination of how variations in one set of parameters affect the overall model, aiding in identifying combinations that yield physically plausible and empirically consistent results.
(a) Cosmological and LF parameters, LSST Y1
(b) Cosmological and LF parameters, LSST Y10
(c) IA and LF parameters, LSST Y1
(d) IA and LF parameters, LSST Y10
(e) LF parameters, LSST Y1
(f) LF parameters, LSST Y10
Figure 9: Segments of the Fisher matrix marginalized contours for JMAS analysis across forecasting years. This figure showcases six plots illustrating the correlations between parameters: the top row details the interactions between cosmological and LF parameters; the middle row explores the correlations between IA and LF parameters; and the bottom row examines the LF parameters’ interrelations. Green contours (left column) represent forecasts for LSST Year 1 (Y1), while blue contours (right column) depict Year 10 (Y10) forecasts, demonstrating the evolution of parameter sensitivities over time.
@@ 5 Summary and Conclusions
  • What are the critical parameter degeneracies of our joint IA-photo-z modelling in the context of a LSST-like weak lensing analysis?
@@
To address these questions § 3 presents a Fisher forecast within the context of both LSST Y1 and Y10, to study the effect of this model extension on cosmological parameter constraints. The forecast is verified to match the modeling guidelines and assumptions defined in the LSST Science Requirements Document (SRD, Mandelbaum et al. 2018). The main differences between the SRD model and ours lies in the parametrization of photometric redshift and IA modelling. The SRD treats uncertainty in the sample redshift distribution by marginalizing over the mean shift and variance of each tomographic source bin. The intrinsic alignment signal is modelled separately. In contrast, we model the luminosity function (LF) separately from the survey volume selection function. We illustrate the impact of the luminosity function parameters on the sample redshift distribution and cosmic shear angular power spectra in § 4.1 and § 4.2. In § 4.3 we marginalize over the luminosity function in the SRD IA model and illustrate the high sensitivity of the Fisher information of the intrinsic alignment model parameters on variations of luminosity function parameters. Specifically, we see that upon marginalizing over the luminosity function parameters the $1\sigma$ constraints degrade by 20%, while the IA parameter constraints are affected up to the 200% level for forecasting year 10. This exploration highlights the fundamental importance and possibilities presented by the complete incorporation of luminosity function parameters into cosmological investigations. By simultaneously modeling systematic influences through a unified luminosity function, we not only capitalize on inherent correlations but also enhance our comprehension of the intricate relationships among different parameters.
In Section § 4.4, we delve into the intricate degeneracies that emerge among cosmological parameters, intrinsic alignment (IA) parameters, and luminosity function (LF) parameters. Notably, we uncover interdependencies within the LF parameter sets and observe their interactions with the broader set of cosmological and IA parameters. This revelation underscores the intrinsic connection between the luminosity function and the cosmic landscape it describes. The interpretation of these findings reveals that such degeneracies are not unexpected, given the physical significance of luminosity function parameters.
Luminosity function parameters, $\phi_{0}^{*}$, $P$, $M_{0}^{*}$, $Q$, and $\alpha$, have a direct impact on the abundance and properties of galaxies as a function of redshift. Consequently, variations in these LF parameters can influence the observed redshift distribution of source galaxies, which in turn affects cosmic shear measurements. This connection between LF parameters and cosmological parameters underscores the value of introducing additional constraints on the luminosity function to disentangle and refine our understanding of these complex relationships. As an illustration of enhancing LF constraints, we use volume selection as an example of a simple selection function for a spectroscopic survey. By incorporating the spectroscopic data, we can enhance the accuracy of cosmological parameter estimation and mitigate potential biases arising from parameter degeneracies.
@@
In our analysis, it is important to note that we do observe a modest decrease in the precision of cosmological parameter constraints when we incorporate the modeling of a luminosity function into our framework. However, it is crucial to emphasize that this reduction in constraints is relatively minor, typically less than a factor of two for most parameters. Importantly, this slight reduction in precision is more than compensated for by the significant advantages gained through a more comprehensive and physically motivated modeling approach.
In fact, our ability to still produce meaningful forecasts for LSST Year 1 and Year 10, despite the slight decline in constraints, underscores the robustness and effectiveness of this modeling technique. The fact that we can provide valuable forecasts for these critical survey years while simultaneously achieving a more consistent and physically grounded model speaks to the power of integrating luminosity function parameters into our cosmological analyses.
@@
Looking ahead, we have plans to further advance this methodology by applying it to Stage III survey data. Additionally, we aim to extend our approach to encompass a joint treatment of sample redshift distribution functions, galaxy-dark matter bias, and intrinsic alignments in the context of a full 3x2pt analysis. These future endeavors highlight the versatility and promise of this modeling technique, which has the potential to significantly enhance our understanding of the universe and its underlying cosmological parameters. In summary, our results demonstrate that this modeling approach is indeed an invaluable asset in the realm of observational cosmology.
@@ Acknowledgements
Software Acknowledgments. The success of this research is largely attributable to several key software tools and packages. Foremost, we acknowledge the foundational role of Python [^2] https://www.python.org. Our analysis heavily relied on essential Python libraries including Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), HDF5 (Group, 2023), and Pandas (pandas development team, 2020). The generation of corner plots was facilitated by the GetDist package (Lewis, 2019), while the CMasher package (van der Velden, 2020) provided invaluable colormaps enhancing our results’ visualization. Finally, the cosmological analysis "heavy lifting" was adeptly supported by the CCL (Chisari et al., 2019), for which we extend our deepest gratitude to the LSST DESC CCL team.
@@ Appendix A Additional corner plots and tables
Figure 10: Comparing SRD and redshift distributions obtained using luminosity functions from GAMA survey. In this figure, we present the LSST Science Requirements Document (SRD) redshift distributions for Y1 and Y10 – represented by green and blue dashed lines, respectively – with distributions derived from a convolution method using luminosity functions (LFs) from the GAMA survey (Loveday et al., 2012). Specifically, we apply the luminosity function parameters for $r$-band red and all galaxy types as detailed in the intrinsic alignment model by Krause et al. 2015. The comparison underscores that the GAMA-based parameters, when used directly, fail to replicate the desired LSST SRD distributions. This discrepancy highlights the necessity to deviate from the GAMA LFs parameters to achieve congruence with the LSST SRD projections.
@@
Figure 11: Comparative analysis between forecasts with (SRD+LF) and without (SRD) luminosity function (LF) parameters. This corner plot illustrates the correlations and distributions of seven cosmological parameters and four intrinsic alignment (IA) parameters within the framework of the LSST cosmic shear analysis. It is divided into two triangular sections, representing forecasts for different years and scenarios. The SRD (Science Requirements Document) scenario, a standard analysis setup (Mandelbaum et al., 2018), utilizes a fixed Schechter luminosity function derived from GAMA and DEEP2 surveys without varying its parameters in IA modeling. In contrast, the SRD+LF (SRD plus Luminosity Function) scenario, as suggested in Krause et al. (2015), expands upon this by varying luminosity function parameters within the IA modeling, providing a more dynamic and potentially more accurate representation of cosmic shear effects. In the lower triangle of the plot, the Year 1 (Y1) forecast is depicted. Here, the SRD+LF scenario is represented by filled chartreuse ellipses, highlighting the parameter space and correlations with varied LF parameters. Overlaid on these are unfilled khaki green contours for the SRD scenario, offering a visual comparative analysis. The filled contours indicate tighter parameter constraints under the SRD+LF scenario, in line with the symmetric percentage differences observed, particularly for parameters like $\Omega_{m}$, $\sigma_{8}$, and the scale factor $n_{s}$ which show modest deviations in Y1. The upper triangle displays the Year 10 (Y10) forecast. Filled neon blue ellipses delineate the SRD+LF scenario, with unfilled denim blue contours for SRD. This section starkly demonstrates the increasing divergence between the two scenarios over time, especially for parameters like $w_{0}$ and $w_{a}$, where the SRD+LF scenario reveals significantly tighter constraints. This is evident in the symmetric percentage differences exceeding $20\%$ for these parameters in the Y10 forecast. Along the diagonal, Gaussian distributions for each parameter indicate the precision of parameter estimates. These distributions underscore the expected uncertainties and central values derived from the Fisher matrices for each scenario and forecasting year. This corner plot not only elucidates the complex correlations between different cosmological and IA parameters but also accentuates the profound influence of integrating luminosity function parameters into IA modeling. The contrasting visualization of SRD and SRD+LF scenarios across Y1 and Y10 forecasts emphasizes the dynamic interplay of parameters and the essential role of comprehensive, adaptive modeling in advancing cosmological understanding through ambitious projects like the LSST.
Figure 12: Relative difference of the joint modelling and standard analysis angular power spectra for forecasting Years 1 and 10. We present the ratios for each correlation, with Year 1 shown in green and Year 10 in blue. The key emphasis here is not the differentiation between individual correlations but to highlight that all observed deviations remain below the 15% threshold. This result is a critical aspect of our validation process, confirming the effectiveness and accuracy of our joint modeling method. By consistently aligning with the SRD benchmarks and maintaining acceptable deviation limits across both forecasting years, this analysis plays a pivotal role in validating our joint modeling approach, showcasing its robustness and reliability in the face of varying forecasting scenarios.

### v3

@@ 4.4.3 Degeneracies: within the luminosity function parameter space
In analyzing the correlations and degeneracies between the Schechter luminosity function parameters themselves, particularly considering their tuning for the LSST Year 1 and Year 10 forecasts, we observe a nuanced interplay. This parameter space is given in Figure. 11. For $\phi_{0}^{*}$, the absence of correlation with $P$ in both Year 1 and Year 10 suggests that the normalization factor of the galaxy number density operates somewhat independently of its evolution parameter. In fact, $\phi_{0}^{*}$ has minimal degeneracy with any of the other luminosity function parameters, which is perhaps unsurprising since it serves only as an overall normalization whereas the others all provide some redshift evolution effects. Parameter $P$ exhibits a strong anti-correlation with $M_{0}^{*}$, $Q$, and $\alpha$ in both years, which can be interpreted as being due to the fact that this set of parameters all induce shifts in the means of the source redshift distributions. The strong positive correlation of $M_{0}^{*}$ with both $Q$ and $\alpha$ in both years, tightening in Year 10, further highlights this effect. $Q$ shows an evolving relationship with $\boldsymbol{\alpha}$, transitioning from a positive correlation in Year 1 to a slightly negative one in Year 10. We expect that this is due to the different fiducial values of the LF parameters, corresponding to a shift in how the rate of change in galaxy luminosity correlates with the abundance of fainter galaxies over time.
The complex interplay between luminosity function (LF) parameters, intrinsic alignment (IA) models, and cosmological parameters underscores the need for an even more extensive investigation. Notably, the parameter space of the LF can be reduced from five to three parameters, as defined in Eq. 4, streamlining our analysis without sacrificing depth or accuracy. Given the intricacies of how these elements influence each other, a comprehensive approach is necessary. A Markov Chain Monte Carlo method, with its capability to dispense with simplifying assumptions of Fisher forecasting (e.g., Gaussian likelihoods), presents a promising route.
@@ 5 Summary and Conclusions
  • What are the critical parameter degeneracies of our joint IA-photo-z modelling in the context of an LSST-like weak lensing analysis?
@@
To address these questions § 3 presents a Fisher forecast set-up within the context of both LSST Y1 and Y10, to study the effect of this model extension on cosmological parameter constraints. We illustrate the impact of the luminosity function parameters on the sample redshift distribution and cosmic shear angular power spectra in § 4.1 and § 4.2. In § 4.3 we marginalize over the luminosity function parameters present in the IA modelling in an otherwise standard analysis approach, and illustrate the high sensitivity of the Fisher information of the intrinsic alignment model parameters on variations of luminosity function parameters. Specifically, we see that upon marginalizing over the luminosity function parameters the $1\sigma$ constraints on dark energy equation of state parameters degrade by 20%, while the IA parameter constraints are affected up to the 200% level for forecasting year 10. This exploration highlights the importance and possibilities presented by the complete incorporation of luminosity function parameters into cosmological investigations.
@@
In Section § 4.4, we compare forecast cosmological results within our full joint analysis framework as compared to a standard analysis approach, and discuss the degeneracies that emerge among cosmological parameters, intrinsic alignment (IA) parameters, and luminosity function (LF) parameters. We find that our approach can produce cosmological parameter constraints which are broadly comparable to that of a standard approach while allowing for greater physical interpretability. We do observe a modest decrease in the precision of some cosmological parameter constraints when we incorporate the modelling of a luminosity function into our framework. However, this slight reduction in precision is arguably compensated for by the significant advantages gained through a more comprehensive and physically motivated modelling approach.
Notably, we uncover interdependencies within the LF parameter sets and observe their interactions with the broader set of cosmological and IA parameters. Luminosity function parameters, $\phi_{0}^{*}$, $P$, $M_{0}^{*}$, $Q$, and $\alpha$, have a direct impact on the abundance and properties of galaxies as a function of redshift. Consequently, variations in these LF parameters can influence the observed redshift distribution of source galaxies, which in turn affects cosmic shear measurements. This connection between LF parameters and cosmological parameters underscores the value of introducing additional constraints on the luminosity function to enhance our understanding of these relationships.
@@
Looking ahead, the framework presented in this work can be extended to encompass a joint treatment of sample redshift distribution functions, galaxy-dark matter bias, and intrinsic alignments in the context of a full 3x2pt analysis. The proof-of-concept offered here also paves the way for application of a similar approach to the treatment of selection function, including with respect to luminosity in redshift, in the more direct estimation of redshift distributions and their joint inference with IA and galaxy bias. These future endeavors highlight the versatility and promise of joint modelling frameworks such as the one presented here, which have the potential to significantly enhance our understanding of the universe and its underlying cosmological parameters.
@@ Acknowledgements
Software Acknowledgments. The success of this research is largely attributable to several key software tools and packages. Foremost, we acknowledge the foundational role of Python [^1] https://www.python.org. Our analysis heavily relied on essential Python libraries including Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), HDF5 (The HDF Group, 2023), and Pandas (The Pandas Development Team, 2020). The generation of corner plots was facilitated by the GetDist package (Lewis, 2019), while the CMasher package (van der Velden, 2020) provided invaluable colormaps enhancing our results’ visualization. Finally, the cosmological analysis "heavy lifting" was adeptly supported by the CCL (Chisari et al., 2019), for which we extend our deepest gratitude to the LSST DESC CCL team.
@@ Appendix A Additional corner plots and tables
Figure 12: Comparing SRD and redshift distributions obtained using luminosity functions from GAMA survey. In this figure, we present the LSST Science Requirements Document (SRD) redshift distributions for Y1 and Y10 – represented by green and blue dashed lines, respectively – with distributions derived from a convolution method using luminosity functions (LFs) from the GAMA survey (Loveday et al., 2012). Specifically, we apply the luminosity function parameters for $r$-band red and all galaxy types as detailed in the intrinsic alignment model by Krause et al. (2016). The comparison underscores that the GAMA-based parameters, when used directly, fail to replicate the desired LSST SRD distributions. This discrepancy highlights the necessity to deviate from the GAMA LFs parameters to achieve congruence with the LSST SRD projections.
@@
Figure 13: Comparative analysis between forecasts with (SRD+LF) and without (SRD) luminosity function (LF) parameters. This corner plot illustrates the correlations and distributions of seven cosmological parameters and four intrinsic alignment (IA) parameters within the framework of the LSST cosmic shear analysis. It is divided into two triangular sections, representing forecasts for different years and scenarios. The SRD (Science Requirements Document) scenario, a standard analysis setup (Mandelbaum et al., 2018), utilizes a fixed Schechter luminosity function derived from GAMA and DEEP2 surveys without varying its parameters in IA modeling. In contrast, the SRD+LF (SRD plus Luminosity Function) scenario, as suggested in Krause et al. (2016), expands upon this by varying luminosity function parameters within the IA modeling, providing a more dynamic and potentially more accurate representation of cosmic shear effects. In the lower triangle of the plot, the Year 1 (Y1) forecast is depicted. Here, the SRD+LF scenario is represented by filled chartreuse ellipses, highlighting the parameter space and correlations with varied LF parameters. Overlaid on these are unfilled khaki green contours for the SRD scenario, offering a visual comparative analysis. The filled contours indicate tighter parameter constraints under the SRD+LF scenario, in line with the symmetric percentage differences observed, particularly for parameters like $\Omega_{m}$, $\sigma_{8}$, and the scale factor $n_{s}$ which show modest deviations in Y1. The upper triangle displays the Year 10 (Y10) forecast. Filled neon blue ellipses delineate the SRD+LF scenario, with unfilled denim blue contours for SRD. This section starkly demonstrates the increasing divergence between the two scenarios over time, especially for parameters like $w_{0}$ and $w_{a}$, where the SRD+LF scenario reveals significantly tighter constraints. This is evident in the symmetric percentage differences exceeding $20\%$ for these parameters in the Y10 forecast. Along the diagonal, Gaussian distributions for each parameter indicate the precision of parameter estimates.